OpenStudy (anonymous):

Simplify showing your steps, tell the property used in each step. 1/2(1*2)+-1

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } (2)+(-1)\] \[\frac{ 1 }{ 2 } (1)\] \[\frac{ 1 }{ 2 } \] I don't understand your question. Are you asking for properties? As in commutative, something like that?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I don't think there are properties to go with those other than closure property. But you see, closure property says that if a and b are real numbers, (a)(b) = unique real number. :)

OpenStudy (anonymous):

ok thank u